# Self-testing in the compiled setting via tilted-CHSH inequalities

### v1

@@ 3.2 Extended definition of the pseudo-expectation
where $\gamma_{i}\in\mathbb{C}$, $k_{i}\in\{0,1\}$ and $w_{i}(B_{0},B_{1})$ is any monomial in the algebra of $\{B_{0},B_{1}\}$. Then there exists a negligible function $\delta_{\mathsf{QHE}}$ of the security parameter $\lambda\in\mathbb{N}$ such that
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[P^{\dagger}P]\geq-\delta_{\mathsf{QHE}}.$ (3.12)
Proof.
To begin, we write
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[P^{\dagger}P]$ $=\sum_{ij}\gamma_{i}^{*}\gamma_{j}\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[(A_{x})^{k_{i}}w_{i}(B_{0},B_{1})(A_{x})^{k_{j}}w_{j}(B_{0},B_{1})]$ (3.13)
$\sum_{ij:k_{i}\oplus k_{j}=0}\gamma_{i}^{*}\gamma_{j}\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\bar{w}_{ij}(B_{0},B_{1})]$ (3.14)
$\approx_{\delta_{\mathsf{QHE}}}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\Bigg{(}\sum_{ij:k_{i}\oplus k_{j}=0}\gamma_{i}^{*}\gamma_{j}\bar{w}_{ij}(B_{0},B_{1})\Bigg{)}|\Psi_{\alpha|\chi}\rangle$
$\sum_{ij}\gamma_{i}^{*}\gamma_{j}\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[(A_{x})^{k_{i}+k_{j}}\bar{w}_{ij}(B_{0},B_{1})]$ (3.15)
$\approx_{\delta_{\mathsf{QHE}}}\sum_{ij:k_{i}\oplus k_{j}=0}\gamma_{i}^{*}\gamma_{j}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\bar{w}_{ij}(B_{0},B_{1})|\Psi_{\alpha|\chi}\rangle$
Where in the fifth line, we used the fact that Bob’s observables satisfy the canonical relations, so we can always replace the canonical monomial $\bar{w}_{ij}$ with $w_{i}w_{j}$. The final line is obtained by noting the square inside the expectation. We therefore conclude $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[P^{\dagger}P]\approx_{\delta_{\mathsf{QHE}}}h$ for some $h\geq 0$, which implies $|\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[P^{\dagger}P]-h|\leq\delta_{\mathsf{QHE}}$, and $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[P^{\dagger}P]\geq h-\delta_{\mathsf{QHE}}\geq-\delta_{\mathsf{QHE}}$ as required. ∎
3.3 Self-testing in the compiled setting
In the compiled setting, a desirable self-testing result would imply that any compiled model achieving the optimal correlation would be efficiently isometric to a canonical compiled model. Furthermore, in cases where the quantum compiled value equals the bipartite value, one would expect isometries relating each compiled model achieving the quantum value to the ideal (bipartite) quantum model.
One of the main obstructions to obtaining strong self-testing results in the compiled setting is the lack of rigidity for the quantum operations applied under the QHE. Nevertheless, in the following we show there are mathematical techniques used to establish self-testing results in the bipartite setting which still provide some partial self-testing results in the compiled setting.
To begin, we recall the description of a bipartite model
$\mathrm{Q}=\big{(}\mathcal{H}_{A},\mathcal{H}_{B},\{\{P_{a|x}\}_{a\in\mathcal{A}}\}_{x\in\mathcal{X}},\{\{Q_{b|y}\}_{b\in\mathcal{B}}\}_{y\in\mathcal{Y}},|\phi\rangle\big{)},$ (3.16)
and consider the set of (sub-normalized) post-measurement states obtained after Alice has measured and $\mathcal{H}_{A}$ is traced out. That is,
$\{\rho_{a|x}:a\in\mathcal{A},x\in\mathcal{X}\}:=\{{\mathrm{t}r}_{A}[(P_{a|x}\otimes\mathbbm{1})|\phi\rangle\langle\phi|]:a\in\mathcal{A},x\in\mathcal{X}\}.$ (3.17)
Definition 3.3.
A bipartite model $\mathrm{Q}$ is said to have a pure $B$-model if $\rho_{a|x}$ is a pure for all $a\in\mathcal{A}$, $x\in\mathcal{X}$. In this case, we write the reduced states as pure states $|\varphi_{a|x}\rangle\in\mathcal{H}_{B}$ satisfying $\rho_{a|x}=|\varphi_{a|x}\rangle\!\langle\varphi_{a|x}|$.
Based on the above, we denote the compiled counter-part of a pure $B$-model $\mathrm{Q}$ as
$\widetilde{\mathrm{Q}}^{*}|_{B}=\big{(}\mathcal{H}_{B},\{|\widetilde{\phi}_{\alpha|\chi}\rangle\}_{\alpha\in\bar{\mathcal{A}},\chi\in\bar{\mathcal{X}}},\{\{Q_{b|y}\}_{b\in\mathcal{B}}\}_{y\in\mathcal{Y}}\big{)},$ (3.18)
where the sub-normalized states $|\widetilde{\phi}_{\alpha|\chi}\rangle$ are given by the relation
$|\widetilde{\phi}_{\alpha|\chi}\rangle:=|\varphi_{a|x}\rangle,\ \forall\chi:\mathsf{Enc}(x)=\chi,\ \forall\alpha:\mathsf{Dec}(\alpha)=a.$ (3.19)
Note that the above implies
$p(a,b,|x,y)$ $=\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha:\mathsf{Dec}(\alpha)=a}\langle\widetilde{\phi}_{\alpha|\chi}|Q_{b|y}|\widetilde{\phi}_{\alpha|\chi}\rangle$ (3.20)
$={\mathrm{t}r}\big{[}(P_{a|x}\otimes Q_{b|y})|\phi\rangle\!\langle\phi|\big{]},$
where we used the inequality ${\mathrm{t}r}_{A}[X_{AB}]Y_{B}={\mathrm{t}r}_{A}[X_{AB}(\mathbbm{1}\otimes Y_{B})]$, telling us that the model $\widetilde{\mathrm{Q}}^{*}|_{B}$ satisfies $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[S]=\eta^{\mathrm{Q}}$.
Based on these definitions, we present a notion of self-testing for compiled correlations. Consider a Bell functional $I$ in the bipartite scenario with Bell operator $S$, a maximum bipartite value $\eta^{\mathrm{Q}}$, and ideal bipartite model $\mathrm{Q}^{*}$. Further suppose that $\mathrm{Q}^{*}$ has compiled counter-part $\widetilde{\mathrm{Q}}^{*}|_{B}$.
Definition 3.4.
The inequality $I\leq\eta^{\mathrm{Q}}$ is a compiled self-test for $\widetilde{\mathrm{Q}}^{*}|_{B}$ if, for every compiled model $\widetilde{\mathrm{Q}}$ which achieves $I\geq\eta^{\mathrm{Q}}-\epsilon$, there exists a negligible function $\delta_{\mathsf{QHE}}$, a function $f(\epsilon)$ such that $f\rightarrow 0$ as $\epsilon\rightarrow 0$, an efficient isometry $V:\widetilde{\mathcal{H}}\mapsto\mathbb{C}^{n}\otimes\mathcal{H}_{\mathrm{aux}}$, and auxiliary states $\{|\mathsf{aux}_{\alpha|\chi}\rangle\}_{\alpha\in\bar{\mathcal{A}},\chi\in\bar{\mathcal{X}}}\subset\mathcal{H}_{\mathrm{aux}}$, which satisfy
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\big{\|}VN_{b|y}|\Psi_{\alpha|\chi}\rangle-Q_{b|y}|\varphi_{\mathsf{Dec}(\alpha)|x}\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\big{\|}^{2}\leq f(\epsilon)+\delta_{\mathsf{QHE}},$ (3.21)
for all $x\in\mathcal{X},b\in\mathcal{B}$, and $y\in\mathcal{Y}$.
Notice that satisfying Eq. 3.21 implies self-testing of the post-measurement states $\{|\Psi_{\alpha|\chi}\rangle\}$ alone. This can be seen by summing both sides over $b\in\mathcal{B}$ and applying the triangle inequality. Although we do not expect to achieve zero error in the compiled setting, we can in theory consider the case where $\epsilon=\delta_{\mathsf{QHE}}=0$. In this case Eq. 3.21 would imply that
$VN_{b|y}|\Psi_{\alpha|\chi}\rangle\!\langle\Psi_{\alpha|\chi}|N_{b|y}V^{\dagger}=Q_{b|y}{\mathrm{t}r}_{A}\big{[}(P_{a|x}\otimes\mathbbm{1}_{B})|\phi\rangle\!\langle\phi|]Q_{b|y}\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\!\langle\mathsf{aux}_{\alpha|\chi}|,$ (3.22)
for all $\alpha$ and $\chi$ such that $a=\mathsf{Dec}(\alpha)$, $\chi=\mathsf{Enc}(x)$. In other words, for all compiled models which maximally violate the Bell expression, we observe that the state after the first round is isometric to preparing a pure bipartite entangled state $|\phi\rangle$ and measuring the first subsystem, according to the honest (ideal) model.
Remark 3.5.
Suppose the ideal nonlocal model employs the maximally entangled state $(|00\rangle+|11\rangle)/\sqrt{2}$, and Alice’s measurements include $A_{0}=\sigma_{Z}$ (as is the case for e.g. the CHSH inequality). Then, in the compiled setting, the honest post-measurement state on Bob’s register after Alice measures $A_{0}$ and gets output $a$ is $|\varphi_{\mathsf{Dec}(\alpha)|0}\rangle=|\mathsf{Dec}(\alpha)\rangle/\sqrt{2}$. Now consider a dishonest device that does not maximally violate the Bell inequality; can we still satisfy Eq. 3.21? One could consider a trivial isometry that ignores the state $|\Psi_{\alpha|\chi}\rangle$ and deterministically prepares $|0\rangle$, via $V|\Psi_{\alpha|\chi}\rangle=(1/\sqrt{2})|0\rangle\otimes|\hat{\Psi}_{\alpha|\chi}\rangle$, where $|\hat{\Psi}_{\alpha|\chi}\rangle=\sqrt{2}|\Psi_{\alpha|\chi}\rangle$ is normalized. Then
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(0)=\chi}\sum_{\alpha}\|V|\Psi_{\alpha|\chi}\rangle-|\varphi_{\mathsf{Dec}(\alpha)|0}\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\|^{2}$ (3.23)
$=\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(0)=\chi}\sum_{\alpha}\frac{1}{2}\||0\rangle-|\mathsf{Dec}(\alpha)\rangle\|^{2}=1.$
Intuitively, any isometry preparing a fixed state independent of $\alpha$ will result in a large distance after summing over the ciphertexts $\alpha$, if the states $|\varphi_{a|x}\rangle$ are orthogonal. This forces the isometry to prepare a state dependent on $\alpha$. Moreover, it must do this according to $\mathsf{Dec}(\alpha)$ to be successful, which would seem to violate the encryption.
4 Tilted-CHSH is a compiled self-test
In this section, we show that the extended tilted-CHSH inequalities satisfy our notion of a compiled self-test.
4.1 Quantum bounds for tilted-CHSH expressions
@@
We now present the family of tilted-CHSH type expressions and their SOS decompositions discovered in [BSB24]. Let $\theta\in(0,\pi/4]$, $\phi\in\big{(}\max\{-2\theta,-\pi+2\theta\},\min\{2\theta,\pi-2\theta\}\big{)}\setminus\{0\}$, and $\lambda_{\theta,\phi}\in\mathbb{R}$ such that
$\frac{1}{\lambda_{\theta,\phi}^{2}}=\frac{\sin^{2}(2\theta)}{\tan^{2}(\phi)}-\cos^{2}(2\theta).$ (4.1)
@@
$S_{\theta,\phi}$ $:=A_{0}\frac{B_{0}+B_{1}}{\cos(\phi)}+\lambda_{\theta,\phi}^{2}\Big{[}\sin(2\theta)A_{1}\frac{B_{0}-B_{1}}{\sin(\phi)}+\cos(2\theta)\frac{B_{0}+B_{1}}{\cos(\phi)}\Big{]},$ (4.2)
$\eta^{\mathrm{Q}}_{\theta,\phi}$ $:=2(1+\lambda_{\theta,\phi}^{2}).$
@@
Let $\theta\in(0,\pi/4]$, $\phi\in\big{(}\max\{-2\theta,-\pi+2\theta\},\min\{2\theta,\pi-2\theta\}\big{)}\setminus\{0\}$, $\lambda_{\theta,\phi}$ be given by Eq. 4.1 and $S_{\theta,\phi},\eta^{\mathrm{Q}}_{\theta,\phi}$ be defined in Eq. 4.2. Define the following polynomials:
$N_{0}$ $:=A_{0}-\frac{B_{0}+B_{1}}{2\cos(\phi)},$ (4.3)
$N_{1}$ $:=A_{1}-\sin(2\theta)\frac{B_{0}-B_{1}}{2\sin(\phi)}-\cos(2\theta)A_{1}\frac{B_{0}+B_{1}}{2\cos(\phi)}.$
The shifted Bell operator $\bar{S}_{\theta,\phi}=\eta^{\mathrm{Q}}_{\theta,\phi}\mathbbm{1}-S_{\theta,\phi}$ admits the SOS decomposition
$\bar{S}_{\theta,\phi}=N_{0}^{\dagger}N_{0}+\lambda_{\theta,\phi}^{2}N_{1}^{\dagger}N_{1}.$ (4.4)
Using the decomposition in Eq. 4.4, it was shown in [BSB24] that the inequality $\langle S_{\theta,\phi}\rangle\leq\eta_{\theta,\phi}^{\mathrm{Q}}$ self-tests the partially entangled state $|\psi_{\theta}\rangle=\cos(\theta)|00\rangle+\sin(\theta)|11\rangle$ and the measurements
$A_{0}$ $=\sigma_{Z},\ A_{1}=\sigma_{X},$ (4.5)
@@
$T_{\theta}=\alpha_{\theta}A_{0}+A_{0}(B_{0}+B_{1})+A_{1}(B_{0}-B_{1}),$ (4.6)
@@
Let $\theta\in(0,\pi/4]$, $\phi\in\big{(}\max\{-2\theta,-\pi+2\theta\},\min\{2\theta,\pi-2\theta\}\big{)}\setminus\{0\}$, and let $S_{\theta,\phi}$ be the extended tilted-CHSH expression with quantum bound $\eta^{\mathrm{Q}}_{\theta,\phi}$, given by Eq. 4.2. Then the maximum quantum value of the corresponding compiled Bell inequality is given by $\eta^{\mathrm{Q}}_{\theta,\phi}+\delta_{\mathsf{QHE}}^{\prime}$, where $\delta_{\mathsf{QHE}}^{\prime}$ is a negligible function of the security parameter.
@@
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\bar{S}_{\theta,\phi}]$ $=\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{0}^{\dagger}N_{0}]+\lambda^{2}_{\theta,\phi}\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{1}^{\dagger}N_{1}],$ (4.7)
@@
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{0}^{\dagger}N_{0}]\geq-\delta_{\mathsf{QHE}}.$ (4.8)
Similarly, $N_{1}$ is expressed in the basis $\{A_{1},B_{0},B_{1}\}$, and we see by Eq. 3.12 that $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{1}^{\dagger}N_{1}]\geq-\delta_{\mathsf{QHE}}$. Putting these together, we obtain
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\bar{S}_{\theta,\phi}]\geq-\delta_{\mathsf{QHE}}(1+\lambda^{2}_{\theta,\phi})=:-\delta_{\mathsf{QHE}}^{\prime},$ (4.9)
which implies $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[S_{\theta,\phi}]\leq\eta^{\mathrm{Q}}_{\theta,\phi}+\delta_{\mathsf{QHE}}^{\prime}$ as desired, where $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[S_{\theta,\phi}]$ is the expected value of the compiled Bell inequality. ∎
@@
The extension of the $S_{\theta,\phi}$ family presented in [[BSB24], Section 3.2.3] self-tests the state $|\psi_{\theta}\rangle$ along with the more general measurements

### v3

@@ 3.2 Extending the pseudo-expectations
where $\gamma_{i}\in\mathbb{C}$, $k_{i}\in\{0,1\}$ and each $w_{i}(B_{0},B_{1})$ is any monomial in the algebra of $\{B_{0},B_{1}\}$. Then there exists a negligible function $\mathrm{negl}(\lambda)$ of the security parameter $\lambda\in\mathbb{N}$ such that
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[P^{\dagger}P]\geq-\mathrm{negl}(\lambda).$ (3.13)
Furthermore, for a given Bell functional $I$, and a compiled model $\widetilde{\mathcal{\mathrm{Q}}}$, $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}(I)$ is the expected value of the compiled model $\widetilde{\mathcal{\mathrm{Q}}}$ on $I$.
The proof can be found in Section A.2.
3.3 Quantum bounds for compiled tilted-CHSH expressions
@@
We now present the family of extended tilted-CHSH type expressions and their SOS decompositions discovered in [BSB24]. Let $\theta\in(0,\pi/4]$, $\phi\in\big(\max\{-2\theta,-\pi+2\theta\},\min\{2\theta,\pi-2\theta\}\big)\setminus\{0\}$, and $\tau_{\theta,\phi}\in\mathbb{R}$ such that
$\frac{1}{\tau_{\theta,\phi}^{2}}=\frac{\sin^{2}(2\theta)}{\tan^{2}(\phi)}-\cos^{2}(2\theta).$ (3.14)
@@
$S_{\theta,\phi}$ $=A_{0}\otimes\frac{B_{0}+B_{1}}{\cos(\phi)}+\tau_{\theta,\phi}^{2}\Big[\sin(2\theta)\,A_{1}\otimes\frac{B_{0}-B_{1}}{\sin(\phi)}+\cos(2\theta)\mathbbm{1}\otimes\frac{B_{0}+B_{1}}{\cos(\phi)}\Big],$ (3.15)
$\eta^{\mathrm{Q}}_{\theta,\phi}$ $=2(1+\tau_{\theta,\phi}^{2}).$
@@
Let $\theta\in(0,\pi/4]$, $\phi\in\big(\max\{-2\theta,-\pi+2\theta\},\min\{2\theta,\pi-2\theta\}\big)\setminus\{0\}$, $\tau_{\theta,\phi}$ be given by Eq. 3.14 and $S_{\theta,\phi},\eta^{\mathrm{Q}}_{\theta,\phi}$ be defined in Eq. 3.15. Define the following polynomials:
$N_{0}$ $=A_{0}\otimes\mathbbm{1}-\mathbbm{1}\otimes\frac{B_{0}+B_{1}}{2\cos(\phi)},$ (3.16)
$N_{1}$ $=A_{1}\otimes\mathbbm{1}-\sin(2\theta)\mathbbm{1}\otimes\frac{B_{0}-B_{1}}{2\sin(\phi)}-\cos(2\theta)\,A_{1}\otimes\frac{B_{0}+B_{1}}{2\cos(\phi)}.$
Then the shifted Bell operator $\bar{S}_{\theta,\phi}=\eta^{\mathrm{Q}}_{\theta,\phi}\mathbbm{1}-S_{\theta,\phi}$ admits the SOS decomposition
$\bar{S}_{\theta,\phi}=N_{0}^{\dagger}N_{0}+\tau_{\theta,\phi}^{2}N_{1}^{\dagger}N_{1}.$ (3.17)
Using the decomposition in Eq. 3.17, it was shown in [BSB24] that the inequality $\langle S_{\theta,\phi}\rangle\leq\eta_{\theta,\phi}^{\mathrm{Q}}$ self-tests following state and measurements:
[…]
$A_{0}$ $=\sigma_{Z},\ A_{1}=\sigma_{X},$
@@
$T_{\theta}=\alpha_{\theta}A_{0}\otimes\mathbbm{1}+A_{0}\otimes(B_{0}+B_{1})+A_{1}\otimes(B_{0}-B_{1}),$ (3.19)
@@
Let $\theta\in(0,\pi/4]$, $\phi\in\big(\max\{-2\theta,-\pi+2\theta\},\min\{2\theta,\pi-2\theta\}\big)\setminus\{0\}$, and let $\langle S_{\theta,\phi}\rangle\leq\eta_{\theta,\phi}^{\mathrm{Q}}$ be the extended tilted-CHSH inequality for the expressions defined in Eq. 3.15. Then the maximum quantum value of the corresponding compiled Bell inequality is given by $\eta^{\mathrm{Q}}_{\theta,\phi}+\mathrm{negl}(\lambda)^{\prime}$, where $\mathrm{negl}(\lambda)^{\prime}$ is a negligible function of the security parameter.
@@
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\bar{S}_{\theta,\phi}]$ $=\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{0}^{\dagger}N_{0}]+\tau^{2}_{\theta,\phi}\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{1}^{\dagger}N_{1}],$ (3.20)
@@
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{0}^{\dagger}N_{0}]\geq-\mathrm{negl}(\lambda).$ (3.21)
Similarly, $N_{1}$ is expressed in the basis $\{A_{1},B_{0},B_{1}\}$, and we see by Eq. 3.13 that $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{1}^{\dagger}N_{1}]\geq-\mathrm{negl}(\lambda)$. Putting these together, we obtain
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\bar{S}_{\theta,\phi}]\geq-\mathrm{negl}(\lambda)(1+\tau^{2}_{\theta,\phi})=:-\mathrm{negl}(\lambda)^{\prime},$ (3.22)
which implies $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[S_{\theta,\phi}]\leq\eta^{\mathrm{Q}}_{\theta,\phi}+\mathrm{negl}(\lambda)^{\prime}$ as desired, where $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[S_{\theta,\phi}]$ is the expected value of the compiled Bell inequality. ∎
@@
The extension of the $S_{\theta,\phi}$ family presented in [BSB24, Section 3.2.3] self-tests the state $|\psi_{\theta}\rangle$ along with the more general measurements